OpenStudy (anonymous):

can someone help me give the number that this expression represents? 10 3/4. 10 1/4 ?

OpenStudy (kira_yamato):

I'll appreciate if you use the equation service provided to illustrate the numbers...

OpenStudy (anonymous):

10&3/4=10.75 10&1/4=10.25

OpenStudy (anonymous):

@kays do u want the answers in decimal fraction or as improper fractions????

OpenStudy (anonymous):

\[\left( \sqrt[5]{8} \right)^{5} \]

OpenStudy (kira_yamato):

8

OpenStudy (anonymous):

ans is 8 as fifth root and raised to power 5 cancel each other out as they r opposite operations

OpenStudy (anonymous):

what is the exponential form for that?

OpenStudy (anonymous):

8^5/5 or 8^1

OpenStudy (anonymous):

THANK YOU , can you help me with the others? i still have , 9 questions here XD pls?

OpenStudy (anonymous):

keep posting them one by one...

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

\[\left( \sqrt[3]{3} \right)^{3}\]

OpenStudy (kira_yamato):

3

OpenStudy (anonymous):

\[\left( \sqrt{36} \right)^{3}\]

OpenStudy (kira_yamato):

6^3 = 216

OpenStudy (anonymous):

can the equation service provided to illustrate the numbers...?

OpenStudy (kira_yamato):

Yeah...

OpenStudy (anonymous):

can you *use

OpenStudy (kira_yamato):

Yeap....

OpenStudy (anonymous):

\[6^{?}\]

OpenStudy (anonymous):

\[(\sqrt{36})^{3} = (6)^3 = 6^{3}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\left( \sqrt[5]{243} \right)^{?}\]

OpenStudy (anonymous):

\[\left( \sqrt[5]{243} \right)^{2}\]

OpenStudy (anonymous):

\[\sqrt[5]{243} = (3^{5})^{1/5} = 3^{5*1/5} = 3^{1}\]

OpenStudy (kira_yamato):

243 = 3^5 \[\sqrt[5]{243}=3\] \[(\sqrt[5]{243})^2=3^2=9\]

OpenStudy (anonymous):

thanks

OpenStudy (kira_yamato):

Anymore?

OpenStudy (anonymous):

yap, wait

OpenStudy (anonymous):

\[(\sqrt[5]{243})^{2} = [(3^5)^1/5]^2 = (3^5/5)^2 = 3^2\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[\left( \sqrt{1/16} \right)^{3}\]

OpenStudy (kira_yamato):

\[=(1/4)^3=1/64\]

OpenStudy (anonymous):

thanks.. \[\left[ 1/2\left( 1/8 \right) \right]^{1/4}\]

OpenStudy (kira_yamato):

\[=\left( \frac{1}{16} \right)^{1/4}=1/2\]

OpenStudy (anonymous):

thanks.. \[\left( 27 2/3 \right)^{-1}\]

OpenStudy (anonymous):

thats 27 and two thirds

OpenStudy (kira_yamato):

\[\left( 27 \frac{2}{3} \right)^{-1}=\left(\frac{83}{3} \right)^{-1}=\frac{3}{83}\]

OpenStudy (anonymous):

thanks.. how do you insert fractions? XD

OpenStudy (kira_yamato):

Just key in frac{"numerator"}{"denominator"}

OpenStudy (anonymous):

thank you, .. \[10 {3}/{4}. 10 {1}/{4}\]

OpenStudy (anonymous):

grrr,, still cant

OpenStudy (anonymous):

thats 10 and 3 fourths thanks :D

OpenStudy (kira_yamato):

\[10 \frac{3}{4} \times 10 \frac{1}{4}?\]

OpenStudy (kira_yamato):

is that what you mean?

OpenStudy (anonymous):

yap :)

OpenStudy (anonymous):

yap

OpenStudy (kira_yamato):

\[10 \frac{3}{4} \times 10 \frac{1}{4} = \frac{43}{4} \times \frac{41}{4} = 110 \frac{3}{16}\]

OpenStudy (anonymous):

thanks \[\left( 25^{2} \right)^{1/4}\]

OpenStudy (kira_yamato):

\[= 25^{\frac{1}{2}} = 5\]

OpenStudy (anonymous):

thanks, last one \[\left( 0.09 \right)^{3/2}\]

OpenStudy (anonymous):

can u give me a medal pls? WHAHAH XD

OpenStudy (kira_yamato):

\[= (0.03)^3 = 2.7 \times 10^{-5}\]

OpenStudy (anonymous):

thanks :D

OpenStudy (kira_yamato):

疲れた。。。

OpenStudy (anonymous):

what do you mean by this 疲れた。。。 ?

OpenStudy (kira_yamato):

Tired.....